Ernie Manufacturing has projected sales of $100.1 million next year. Costs are expected to be $90 million and net investment is expected to be $5 million. Each of these values is expected to grow at 14 percent the following year, with the growth rate declining by 2 percent per year until the growth rate reaches 6 percent where it will remain. There are 5.5 million shares of stock outstanding. Investors require a return of 13 percent and the corporate tax rate is 40 percent.

What is your estimate of the current stock price?
Answer #1

SOLUTION -

Calculation of free cash flow at year 1

  • Projected Sales = $100.1 million
  • Projected Cost = $90 million
  • Projected Income before Tax = $100.1 million - $90 million = $10.1 million
  • Projected Income after Tax = $10.1 million*(1-0.4) =$6.06 million
  • Projected Free Cash flow = $6.06 million - $5 million = $1.06 million

Calculation of growth for Free Cash Flow from year 1:

Year Growth rate Value (IN $)
1 ..... 1060000
2 14% 1208400
3 12% 1353408
4 10% 1488749
5 8% 1607849
6 6% 1704320
7 6% 1806579
n 6% ......

a. Calculation of Value

We calculate the firm value by considering the present value of free cash flows

  • Step 1: we calculate terminal value at year 5 (after year 5 the growth rate became stable at 6%)

TV(5) = CF(6)/(k-g) = $1704320/(0.13/0.06) = $24347429

  • Step 2: We calculate the present value of all cash flows (from year 1 to year 5) and the terminal value

PV=$1060000/1.13+$1208400/(1.13)^2+$1353408/(1.13)^3+$1488749/(1.13)^4+$1607849/(1.13)^5+$24347429/(1.13)^5....................(1)

    • = $178,20,721
  • Step 3:

Valuation of each Share = Total Value/Total Number of shares = $178,20,721/5.5 million

= $3.24 per Share.